# Need Help? Designing Proactive AI Assistants for Programming

### v1

@@ 4 Proactive Assistant Implementation
In this paper, we build a proactive assistant for programming tasks based on the design considerations in Section 3. Our proactive chat assistant distinguishes itself from online AI assistants such as ChatGPT in two ways: first, our proactive assistant has access to the user’s work context, and second, our proactive assistant proposes suggestions instead of passively waiting for user requests. When embedding the proactive assistant into the user’s work context—i.e., the programmer’s IDE, the proactive assistant will have access to both the current code in any of the user’s files and prior user queries to the AI assistant. To turn context into suggestions, we discuss two primary components that comprise our proactive assistant, as shown in Figure 1, and how they instantiate the set of design considerations:
@@
  • Interactive interface: In Section 4.1, we overview the chat interface which shows suggestions from the proactive assistant in a way that is easy for users to interact with and understand (Design Considerations 1-2) and to provide feedback (Design Consideration 5).
@@
  • Suggestion generation: In Section 4.2, we discuss how the assistant takes the current context and generates a set of proactive suggestions that is timely and relevant (Design Consideration 3). In Section 4.3, we discuss how to appropriately determine when suggestions are shown to users (Design Consideration 4).
4.1 Interactive Interface
@@
Figure 2: A walkthrough of the proactive assistant interface for coding. (A) Overview of suggestions, where the assistant provides a short description, and users can choose to expand a suggestion for more details (this can involve implementation and a brief explanation); (B) Buttons that allow participants to preview the implementation in their code, accept a suggestion to ask follow-up questions on, or delete the suggestion; (C) Integrating a suggestion into the editor via a diff format, where users can decide if they want to accept or hide the changes; (D) Invoking suggestions, where users can also request suggestions from the assistant (for example, if they did not like the original suggestion or wanted a suggestion before the assistant provided one).
@@
The interface of the proactive assistant, shown in Figure 2, is built into the standard chat interface, allowing programmers to use all normal chat functionality, while the assistant periodically provides suggestions to programmers. We describe the different features of the proactive suggestion and their respective design considerations:
4.1.1 Suggestion summary.
@@
Towards fulfilling design consideration 1, the proactive assistant facilitates efficient evaluation of the suggestion by providing a summary to allow users to decide whether the suggestion is relevant. The summary consists of a single sentence that starts with the type of the suggestion e.g., a bug fix, a new feature, and then a description of the suggestion, as shown in the headings of Figure 2 (A). The drop-down style of the interface allows users to browse the various suggestions easily. Through pilot studies, we find that showing a set of three suggestions strikes the balance between creating a diverse yet manageable set of information for users to look at. We explore the effect of changing this parameter in our user study.
4.1.2 Details of suggested implementation.
@@
If the user decides to look into a suggestion further they can expand it for more details as shown in Figure 2 (A). Depending on the type of suggestion, they may receive a code snippet and/or a text description (e.g., a suggestion that explains a code snippet may not include code). The interface also allows users to copy the suggestion into the editor (design consideration 2). We limit the text description of the implementation to a few bullet points to facilitate efficient interaction (design consideration 1).
4.1.3 Preview suggested implementation in code.
@@
While users can copy the suggestion and decide how to use it themselves, the proactive assistant also has a “preview” functionality where the assistant will suggest an option for how the suggestion can be incorporated into the code (design consideration 2). The suggested implementation is computed via another call to the LLM, where it is given the proactive suggestion and the user’s code and produces new code incorporating the suggestion. This generated code file is then shown to the user via a diff editor based on the Monaco Editor [^1] https://microsoft.github.io/monaco-editor/ which tells the user what lines were added or removed once the suggestion was incorporated, as shown in Figure 2 (C). Once presented with the preview, the user decides whether to accept or hide the changes. They also accept a subset of the changes by editing freely before clicking on the “accept changes” button. In the Appendix, we include the prompt that incorporates edits into the current code snippet.
4.1.4 Asking follow-up questions based on suggestion.
@@
A natural interaction paradigm when using chat assistants is for users to provide clarifications or ask follow-up questions based on the assistant responses. Since the proactive assistant is built into the chat interface, we tap into this paradigm when scaffolding interactions for proactive suggestions. As such, we create the option for users to add a proactive suggestion to the chat history (via the “accept” button) or remove it (via the “delete” button). Clicking on the accept button shown in Figure 2 (B) would add the proactive suggestion to the chat message in a way that looks visually similar to the rest of the user’s messages. This also allows the user to ask follow-up questions on the proactive suggestion. This may be helpful as it is common for users to follow up on LLM responses to clarify or correct the model. Further, accepting and deleting suggestions allows future proactive suggestions to incorporate user feedback, addressing design consideration 5.
4.1.5 Manually requesting suggestions.
@@
While the nature of a proactive assistant suggests that the assistant will determine when to provide suggestions to the user, it is unrealistic to expect an assistant to anticipate when users might want suggestions (design consideration 4). As such, we include a button, as shown in Figure 2 (D), to allow users to request suggestions. This can also allow us, in future work, to better learn when the proactive assistant should intervene.
4.2 Suggestion Content
@@
Completing unfinished code The following is some C code for binding a "hello" C function to an Io method. Can you complete the code by the comment:
@@
Table 1: Suggestion types informed by dataset analysis. We identify categories of suggestions that a proactive assistant can surface by coding developer questions to ChatGPT in the DevGPT [6] dataset. We incorporate these suggestion types into the design of our proactive assistant.
@@
For the proactive assistant to be useful, it must generate appropriate suggestion content (design consideration 3). This means that the assistant should understand what types of suggestions will benefit users and that the assistant should surface the right type of suggestion at the right time, depending on the user’s code context.
@@
To identify the different types of suggestions that users might ask and benefit from, we analyze DevGPT [6], a dataset of conversations between programmers and ChatGPT, as a source of real-world questions. While other chat datasets also contain in-the-wild programming-related questions (e.g., WildChat [50], LMSYS-Chat-1M [51]), we selected DevGPT in particular because the majority or all questions contain user code context in addition to the user question and LLM response, which allows us to manually inspect whether it would be reasonable for a proactive assistant to anticipate this question from the current code context. We sample 100 questions and their code contexts from the DevGPT data and manually create a label for the “category” of the question. We then cluster labels to identify common themes across questions. The goal of surveying existing questions is to ensure that our proactive assistant can provide diverse suggestions; however, the categories we identify may not be exhaustive.
@@
From this annotation process, we identify eight categories of questions that proactive assistants can help with: explaining existing code, brainstorming new ideas or functionality, completing unfinished code, providing pointers to syntax hints or external documentation, identifying and fixing bugs (which include both latent and runtime errors), adding unit tests, and improving code efficiency and modularity. In Table 1, we include example user questions that fall under each category. To scaffold these suggestion types into proactive suggestions, we incorporate them into the system prompt to encourage the model to consider a more diverse set of suggestions. In the Appendix, we include the prompt used to generate the proactive suggestions.
4.3 Timing of Suggestions
@@
The proactive assistant must time suggestions appropriately to not be obstructive or distracting to users (design consideration 4). A study by Barke et al. [20] characterized interactions with programming assistants as consisting of two modes: acceleration, where the assistant is primarily used to help the programmer implement what they already intend to write; and exploration, where the programming assistant assists a programmer in identifying and planning out goals. We propose the following conditions for when proactive suggestions should be shown to users based on an estimation of when the user is in acceleration or exploration mode:
@@
  • During acceleration: Suggestions are not requested when the user is interacting with a previous suggestion, typing in the chat, sending a message, or waiting for a response. The suggestion timer resumes 5 seconds after the user stops typing to account for a typing break. If the user would like a suggestion, they can manually request one.
@@
  • During exploration: When the user is idle—likely due to planning and brainstorming, the assistant provides suggestions every 5 seconds, limited to every 20 seconds since the last suggestion, interaction, or chat. If the user starts coding before suggestions are shown, due to potential latency in the model query, we do not display the suggestions.
@@
Additionally, we show a proactive suggestion during debugging, when users are trying to edit the current code for any bugs or improve the performance. Debugging is distinct from the previous two states because the user is neither trying to write more code nor planning the next steps. The proactive assistant is immediately triggered when the user runs or submits code that leads to an error because the user is not actively coding at the moment. We modify the suggestion generation prompt to incorporate the error statement and restrict the types of suggestions that the assistant can generate. In the Appendix, we include the prompt used to generate suggestions during debugging.
5 Study Design
@@
    What is the effect of the proactive assistant on user productivity? If the proactive chat assistant can anticipate user needs or even surface issues the user has not considered, we expect the assistant to have a measurable downstream effect on productivity measures.
@@
    What is the effect of the proactive assistant on user experience? In addition to productivity measures, the user should have a favorable experience when interacting with proactive assistants. If the proactive assistant is providing sufficient perceived value-add, the user would ideally perceive the proactive assistant to be more helpful than a standard assistant like ChatGPT.
@@
5.1 Participants
We recruited a total of 61 students via university mailing lists. The inclusion criteria for the study are that they must be based in the U.S., be older than 18 years old, and have experience programming. Since we design tasks that are written in Python, we require participants to have a baseline level of Python knowledge. Among our participants, 60% reported their Python knowledge as intermediate, 31% had advanced Python experience, and 9% were beginners. Moreover, 19% self-report that they are daily users of AI tools for programming (including GitHub Copilot and ChatGPT), 46% report that they use these tools at least once if not multiple times a week, and 35% rarely use AI tools for programming.
5.2 Baseline
Our study compares proactive assistants to a non-proactive chat-based assistant baseline. Every participant in our study interacts with a standard chat assistant that is incorporated into the coding environment but does not have access to the user’s code. This is similar to the set-up that has been studied in prior papers [31, 6, 18] and mimics how people would use ChatGPT. The interaction paradigm for the baseline is as follows: after the user types in their query, which may contain code from the IDE to provide relevant context, the chat assistant can respond in natural language and provide code snippets. To utilize the code snippets, the user would need to copy in the assistant response back into the IDE.
In both the baseline and our experimental conditions, we do not include any code autocomplete tools to not introduce a potential confounding factor in our analysis as not all participants are experienced in using autocomplete tools. Further, we believe the proactive assistant can complement code autocompletion tools, which consist of local proactive changes, while the proactive assistant provides suggestions at a more global level.
5.3 Experimental Conditions
@@
In this condition, the participant interacts with the proactive assistant described in Section 4, without the preview feature that allows the assistant to suggest direct edits to the user’s code. Similar to Condition 1, the participant can still send chat messages to the assistant at any point while also receiving proactive suggestions.
@@
In this condition, the proactive assistant in this condition looks at face value to be the same as the proactive assistant in Condition 2 (Suggest). However, we modified a few different parameters that were discussed in Section 4: we reduced the amount of time that an assistant waits to provide a suggestion (from 20 seconds to 5 seconds), increased the number of suggestions shown (from 3 to 5), and removing guiding prompts.
@@
The participants were randomized and roughly divided equally across the three conditions, with 20 in Suggest and Preview, 21 in Suggest, and 20 in Persistent Suggest. Comparing the three conditions allows us to investigate whether (1) the level of proactivity affects user productivity and user experience (comparing Suggest and Preview and Suggest) and whether (2) small changes in the design of a proactive assistant can make a sizeable difference in user productivity and experience (comparing Suggest and Persistent Suggest). We fix the LLM backbone across all conditions to be GPT-4o [1], a state-of-the-art LLM.
5.4 Experimental Platform

### v2

@@ 4. Proactive Assistant Implementation
In this paper, we build a proactive assistant for programming tasks based on the design considerations in Section 3. Our proactive chat assistant distinguishes itself from online AI assistants such as ChatGPT by first having access to the user’s work context and second by proactively proposing suggestions to the user versus only passively waiting for user requests. When embedding the proactive assistant into the user’s work context, which is the programmer’s IDE in this setting, the proactive assistant will have access to the current code in any of the user’s files, the terminal outputs, and prior user queries to the AI assistant. To turn the context into suggestions, we discuss different components that comprise our proactive assistant and how we instantiate each design consideration, as shown in Figure 1:
@@
  • Interactive interface: In Section 4.1, we overview the interface through which the assistant presents the generated suggestions in a way that makes it easy for users to interact with, understand, and invoke when needed (Design Considerations 1-2).
@@
  • Suggestion generation: In Section 4.2, we discuss how the assistant takes the current context and generates a set of proactive suggestions that is timely and relevant (Design Consideration 3) which take into account user feedback (Design Consideration 4). In Section 4.3, we discuss how to appropriately determine when suggestions are shown to users (Design Consideration 5).
4.1. Interactive Interface
@@
Figure 2. A walkthrough of the proactive assistant interface for coding. (A) Overview of suggestions, where the assistant provides a short description, and users can choose to expand a suggestion for more details (this can involve implementation and a brief explanation); (B) Buttons that allow participants to preview the implementation in their code, accept a suggestion to ask follow-up questions on, or delete the suggestion; (C) Integrating a suggestion into the editor via a diff format, where users can decide if they want to accept or hide the changes; (D) Invoking suggestions, where users can also request suggestions from the assistant (for example, if they did not like the original suggestion or wanted a suggestion before the assistant provided one).
[…]
A walkthrough of the proactive assistant interface for coding. (A) Overview of suggestions, where the assistant provides a short description, and users can choose to expand a suggestion for more details (this can involve implementation and a brief explanation); (B) Buttons that allow participants to preview the implementation in their code, accept a suggestion to ask follow-up questions on, or delete the suggestion; (C) Integrating a suggestion into the editor via a diff format, where users can decide if they want to accept or hide the changes; (D) Invoking suggestions, where users can also request suggestions from the assistant (for example, if they did not like the original suggestion or wanted a suggestion before the assistant provided one).
@@
The interface of the proactive assistant, shown in Figure 2, is built into the standard chat interface, allowing programmers to use all normal chat functionality, while the assistant periodically provides suggestions to programmers. Furthermore, this means the proactive assistant does not take up additional real estate on the screen, particularly when programmers may already have multiple files open side-by-side. We describe the different features of the proactive suggestion and their respective design considerations:
4.1.1. Suggestion summary.
@@
Towards fulfilling design consideration 1, the proactive assistant facilitates efficient evaluation of the suggestion by providing a summary to allow users to quickly get an idea of whether or not the suggestion is relevant. The summary consists of a single sentence that starts with the type of the suggestion e.g., a bug fix, a new feature, and then a description of the suggestion, as shown in the headings of Figure 2 (A). The drop-down style of the interface allows users to easily expand and condense a suggestion and quickly browse the various suggestions.
4.1.2. Details of suggested implementation.
@@
If the summary of the suggestion seems relevant to the user, they can expand the suggestion for more details as shown in Figure 2 (A). Depending on the type of suggestion, they may receive a code snippet and/or a text description (e.g., a suggestion that explains a code snippet may not include code). The interface also allows users to copy the suggestion into the editor (design consideration 2). We limit the text description of the implementation to a few bullet points to facilitate efficient interaction (design consideration 1).
4.1.3. Preview suggested implementation in code.
@@
While users have the freedom to copy the suggestion code into the editor and decide how they want to use the suggestion, the proactive assistant also has a “preview” functionality, which will allow users to see how the assistant would incorporate it into their code (design consideration 2). The suggested implementation is computed via another call to the LLM, where it is given the proactive suggestion and the user’s code and produces a new code that incorporates the suggestion. Each time a suggestion is previewed requires another LLM inference call. As such, we only preview a suggestion when invoked by the user, rather than automatically previewing every suggestion, which helps to reduce the latency that users experience. This new code file generated is then shown to the user via a diff editor based on the Monaco Editor  [^1] https://microsoft.github.io/monaco-editor/ which tells the user what lines were added or removed once the suggestion was incorporated, as shown in Figure 2 (C). The user has the option to accept or revert all changes and can additionally only a subset of the changes and edit freely before clicking on the “accept changes” button. In the Supplementary Material, we include the prompt used to incorporate edits into the current code snippet.
4.1.4. Asking follow-up questions based on suggestion.
@@
Since the proactive assistant is a part of the chat interface, we create the option for users to add a proactive suggestion to the chat history (via the “accept” button) or remove it (via the “delete” button). Clicking on the accept button shown in Figure 2 (B) would add the expanded suggestion to the chat message in a way that looks visually similar to the rest of the user’s messages. This allows future proactive suggestions to condition on the user’s accept or delete actions addressing design consideration 4 (incorporating user feedback). This also allows the user to conveniently ask follow-up questions on the proactive suggestion. This may be helpful as it is common for users to follow up on LLM responses to clarify or correct the model.
4.1.5. Manually requesting suggestions.
@@
While the nature of a proactive assistant suggests that the assistant will be the one who determines when to provide suggestions to the user, it is unrealistic to expect an assistant always to anticipate when users might want suggestions (design consideration 5). As such, we include a button, as shown in Figure 2 (D), allowing users to request suggestions as they wish. This can also allow us in future work to better learn when the proactive assistant should intervene.
4.2. Suggestion Content
@@
Completing unfinished code The following is some C code for binding a ”hello” C function to an Io method. Can you complete the code by the comment:
@@
Table 1. Suggestion types informed by dataset analysis. We identify categories of suggestions that a proactive assistant can surface by coding developer questions to ChatGPT in the DevGPT (Xiao et al., 2023) dataset. We incorporate these suggestion types into the design of our proactive assistant.
[…]
Suggestion types informed by dataset analysis. We identify categories of suggestions that a proactive assistant can surface by coding developer questions to ChatGPT in the DevGPT (Xiao et al., 2023) dataset. We incorporate these suggestion types into the design of our proactive assistant.
@@
For the proactive assistant to be useful, it must generate appropriate suggestion content (design consideration 3). This means that the assistant should have knowledge of the types of suggestions that users benefit from and that the assistant should surface the right type of suggestion at the right time, depending on the user’s code context.
@@
To identify the different types of suggestions that users benefit from for coding contexts, we look to DevGPT (Xiao et al., 2023), a dataset of conversations between programmers and ChatGPT, as a source of real-world questions. While other chat datasets also contain in-the-wild programming-related questions (e.g., WildChat (Zhao et al., 2024), LMSYS-Chat-1M (Zheng et al., 2023)), we selected DevGPT in particular because the majority or all questions contain user code context in addition to the user question and LLM response, which allows us to manually inspect whether we think it would be reasonable for a proactive assistant to anticipate this question from the current code context. We sample 100 questions and their code contexts from the DevGPT data and manually create a label for the “category” of the question. We then cluster labels to identify common themes across questions. The goal of surveying existing questions is to ensure we obtain a reasonable set of potential suggestion types to prompt the proactive assistant, however, this set may not be exhaustive.
@@
From this annotation process, we identify eight categories of questions that proactive assistants can help with: explaining existing code, brainstorming new ideas or functionality, completing unfinished code, providing pointers to syntax hints or external documentation, identifying and fixing bugs (which include both latent and runtime errors), adding unit tests, and improving code efficiency and modularity. Annotations were roughly distributed across all categories where each appeared 15%, 19%, 18%, 13%, 10%, 9%, 6%, 10% of the time respectively—note that since this dataset consists of only ChatGPT conversations and may not represent the actual distribution of user questions in practice. In Table 1, we include example user questions that fall under each category.
4.3. Timing of Suggestions
@@
The proactive assistant must display suggestions at the appropriate time so they are not obstructive and distracting to users (design consideration 5). A study by Barke et al. (2023) characterized that interaction with programming assistants consists of two modes: acceleration, where the assistant is primarily used to help the programmer implement what they already intend to write; and exploration, where the programming assistant assists a programmer in identifying and planning out goals. We propose the following conditions for when proactive suggestions should be shown to users based on an estimation of when the user is in acceleration or exploration mode:
@@
  • During acceleration: Suggestions are not requested when the user is interacting with a previous suggestion, typing in the chat, sending a message, or waiting for a response. The suggestion timer resumes 5 seconds after the user stops typing to account for a typing break. If the user would like a suggestion, they can still request it.
@@
  • During exploration: When the user is idle—likely due to planning and brainstorming, the assistant provides suggestions after 5 seconds, limited to every 20 seconds since the last suggestion, interaction, or chat. If the user starts coding before suggestions are shown, due to potential latency in the model query, we do not display the suggestions.
To detect mode based on user interactions, we implement listeners and corresponding timers in the editor to track whether the user is actively taking actions (e.g., writing code or chat messages). Based on pilot studies with 4 participants, we determined that 5 seconds was a reasonable estimates of how long people tend to pause when coding and 20 seconds was a reasonable guess at how much time was needed to check out the proactive suggestions.
@@
Additionally, we allow the proactive suggestion to be shown during debugging when users are trying to edit the current code for any bugs or improve the performance. Debugging is distinct from the previous two states because the user is neither trying to write more code nor planning the next steps. The proactive assistant immediately provides a suggestion when the user runs or submits code that leads to an error because the user is not actively coding.
Figure 3. Prompting the assistant for suggestions. An overview of how different inputs (e.g., prior messages, current code, and terminal outputs) are incorporated in the generation of standard and debugging suggestions, the former which is shown generally and the latter which is only triggered when code is run. Full prompts for both types of suggestions are provided in the Supplementary Material.
[…]
Prompting the assistant for suggestions. An overview of how different inputs (e.g., prior messages, current code, and terminal outputs) are incorporated in the generation of standard and debugging suggestions, the former which is shown generally and the latter which is only triggered when code is run. Full prompts for both types of suggestions are provided in the Supplementary Material.
4.4. Generating Suggestions
To generate a set of suggestions, we discuss how the assistant makes a call to the LLM with the current user context to generate a set of suggestions. In this work, all LLM calls are made to GPT-4o (OpenAI, 2022), a state-of-the-art LLM. In pilot studies, we experimented with smaller, faster models (e.g., GPT-4o-mini) but observed that the quality (e.g., correctness of suggested code, relevance of suggestions) of suggestions to be lower. We overview what inputs from the user context are used to generate each type of suggestion and include the full prompts in the Supplementary Material.
Standard suggestions.
Standard suggestions are the ones that periodically show up while the user is writing code. To generate these suggestions, the LLM call includes a list of items: any prior chat messages, a system prompt, and a general prompt. We include the prior messages to avoid generating suggestions that may repeat previously asked questions and to make the model aware of prior user-assistant interactions. Then, we scaffold the suggestion types identified in Section 4.2 into the system prompt to encourage the model to consider a more diverse set of suggestions. Finally, we include the current code and instructions for how to return the set of generated suggestions in the main prompt. Figure 3 (left) shows how all inputs are combined to generate standard suggestions.
Debugging suggestions.
Debugging suggestions are the suggestions that are only triggered when the user runs their code. The LLM call again includes prior messages, but instead of the system prompt, we simply include all other information in one main prompt. This main prompt includes the current code, a short instruction to ask the model to focus on any error outputs, error statements from the terminal, and instructions for how to return the set of generated suggestions. Note that debugging suggestions have a more restricted set of suggestions that the assistant can generate. Figure 3 (right) shows how all inputs are combined to generate debugging suggestions.
The response from the LLM will then be parsed into individual suggestions and displayed in the exact order in which they are returned in the interactive interface as described in Section 4.1. While we do not explicitly incorporate an additional mechanism to rank suggestions, we verify based through pilot tests that the order of suggestion types is not always the same.
5. Study Design
@@
    What is the effect of the proactive assistant on user productivity? The goal of a proactive chat assistant is to anticipate potential user needs or even surface questions the user has not considered. As such, we expect the assistant to have a measurable downstream effect on productivity measures.
@@
    What is the effect of the proactive assistant on user experience? In addition to productivity measures, we hope the user has a favorable experience when interacting with proactive assistants.
@@
5.1. Participants
We recruited a total of 65 students via university mailing lists. The inclusion criteria for the study are that they must be based in the U.S., be older than 18 years old, and have experience programming. Since we design tasks that are written in Python, we require participants to have a baseline level of Python knowledge. Among our participants, 60% reported their Python knowledge as intermediate, 31% had advanced Python experience, and 9% were beginners. Moreover, 19% self-report that they are daily users of AI tools for programming (including GitHub Copilot and ChatGPT), 46% report that they use these tools at least once if not multiple times a week, and 35% rarely use AI tools for programming. The gender breakdown of our participants was 34% female and 66% male. The student population was distributed between 38% undergraduate and 62% graduate students, skewing towards more experienced programmers.
5.2. Baseline
Our study compares proactive assistants to a non-proactive chat-based assistant baseline. Every participant in our study interacts with a standard chat assistant that is incorporated into the coding environment but does not have access to the user’s code. This is similar to the set-up that has been studied in prior papers (Kazemitabaar et al., 2024; Xiao et al., 2023; Mozannar et al., 2024c) and mimics how people would use ChatGPT. The interaction paradigm for the baseline is as follows: the user needs to type in their query to obtain a response from the chat assistant. The chat assistant can respond in natural language and provide code snippets.
In the baseline (as well as our experimental conditions), we do not include any code autocomplete tools to not introduce a potential confounding factor in our analysis as not all participants are experienced in using autocomplete tools. Further, we believe the proactive assistant can complement code autocompletion tools, which consist of local proactive changes, while the proactive assistant provides suggestions at a more global level.
5.3. Experimental Conditions
@@
In this condition, the participant interacts with the proactive assistant described in Section 4, without the preview feature, which allows the assistant to suggest direct edits to the user’s code. Similar to Condition 1, the participant can still send chat messages to the assistant at any point while also receiving proactive suggestions.
@@
In this condition, the participant interacts with a variant of our proactive chat assistant. While the proactive assistant in this condition looks at face value to be the same as the proactive assistant in Condition 2 (Suggest), we modified a few different parameters discussed in Section 4: reducing the amount of time that an assistant waits to provide a suggestion (from 20 seconds to 5 seconds), increasing the number of suggestions shown (from 3 to 5), and removing guiding prompts.
@@
The three conditions allow us to investigate whether (1) the level of proactivity affects user productivity and user experience (comparing Suggest and Preview and Suggest) and whether (2) small changes in the design of a proactive assistant can make a sizeable difference in user productivity and experience (comparing Suggest and Persistent Suggest). Recall we fix the LLM backbone across all conditions to be GPT-4o (OpenAI, 2022), a state-of-the-art LLM.
5.4. Experimental Platform